OpenStudy (anonymous):

Help!!! 2. The perimeter of a rectangular flower garden is 60 m and its area is 225 m2. Find the length of the garden.

OpenStudy (anonymous):

 Based on this drawing of the rectangular flower garden, the area of the garden will by X*Y, and the perimeter will be Y+X+Y+X, or P = 2X+2Y. So now you have: A = X*Y P = 2X+2Y if you solve the first equation to give X in terms of Y, you get X = A*Y You then plug that into the perimeter equation to get: P = 2*(X) + 2*Y or, P = 2*(A*Y) + 2*Y From there, you can solve for Y using your given A and P, and then solve for X using Y